OpenStudy (anonymous):

How do you graph this parametric equation? x = 2t y = t + 5, -2 ≤ t ≤ 3

OpenStudy (anonymous):

@jim_thompson5910 Do you have any ideas?

jimthompson5910 (jim_thompson5910):

when t = -2, what is x? what is y?

jimthompson5910 (jim_thompson5910):

alternatively, you can solve y = t + 5 for t to get y = t + 5 y-5 = t t = y-5 Then plug that into x = 2t to get x = 2t x = 2(y-5) I'll let you isolate y